OpenStudy (anonymous):

Given the vectors u=(-2,3) and v=(2,1), determine a vector w as 2(u-w)+4(v+w)=0

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

are you sure you copy the right problem? just make sure about what we should do

OpenStudy (loser66):

@Teuzer knock, knock

OpenStudy (anonymous):

ok

OpenStudy (loser66):

let w (x,y ) so tell me what u-w =??

OpenStudy (anonymous):

in this problem 0= zero vector

OpenStudy (loser66):

let it aside, not touch it yet

OpenStudy (loser66):

just answer me, u-w =...

OpenStudy (anonymous):

don't know

OpenStudy (loser66):

my god, u -w = (-2 -x, 3 -y) . you really not know about that? are you kidding me?

OpenStudy (anonymous):

ohh. I know this. but how can I discovery w ?

OpenStudy (loser66):

take steps, please.!! ok, so how about 2 (u - w) =?

OpenStudy (anonymous):

I believe the final result is w=(2,5)

OpenStudy (anonymous):

Is that right?

OpenStudy (loser66):

how do you get it? teach me back, please

OpenStudy (loser66):

@phi rescue, please

OpenStudy (anonymous):

wel. 2(-2-x,3-y) + 4(2+y,1+y) (-4-2x,6-2y)+(8+4y,4+4y) (4+2x,10+2y) x=4/2 and y= 10/2

OpenStudy (loser66):

the last part is wrong, it is 4 ( 2 +x , 1 +y ) not 4 (2 +y ,....) that's mistake,

OpenStudy (loser66):

you redo, that's the right set up , just fix that mistake . show me your work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I just wrote it wrong. but the result is right

OpenStudy (loser66):

nope, re calculate

OpenStudy (anonymous):

so wherer it's wrong? i don't see it.

OpenStudy (loser66):

first term for x: -4 -2x +8 +4x = 4 +2x = 0 ---> 2x = -4---> x =-2

OpenStudy (loser66):

see it?

OpenStudy (anonymous):

Ohh yeah. just the signal.

OpenStudy (loser66):

thnks phi for reply my tag. :)

OpenStudy (loser66):

value of y is the same. @Teuzer it not "just the signal" in linear, it's wrong.

OpenStudy (anonymous):

I got, i meant my mistake was upon the signal.

OpenStudy (phi):

***Given the vectors u=(-2,3) and v=(2,1), determine a vector w as 2(u-w)+4(v+w)=0 I would divide both sides by 2 to simplify 2(u-w)+4(v+w)=0 u - w + 2(v+w)=0 distribute the 2 to get u - w +2v + 2w= 0 combine 2w-w to get w: u + 2v + w = 0 add -u -2v to both sides w = -u -2v notice this is like ordinary algebra. now replace u and v with their vector values. combine like components.

OpenStudy (loser66):

hehehee.... that's why I tag him, thanks phi . Perfect performance.

OpenStudy (phi):

You still have to find the vector w.

OpenStudy (anonymous):

Good thinking.

OpenStudy (loser66):

@phi, the leftover is the Asker's responsibility. easiest way, i gave out, professional way, you gave out. that's it.

OpenStudy (phi):

to finish this, what is -u if u is (-2,3) ? what is -2v if v is (2,1) ? once you get those vectors add corresponding components.